# Automated Genomic Interpretation via Concept Bottleneck Models for Medical Robotics

### v1

@@ I Introduction
As medical robots are deployed for triage, specimen handling, and intra-/post-operative decision support, their perception and decision modules largely rely on end-to-end deep learning, whose “black-box” nature undermines trust in safety-critical contexts. Recent work shows that explainable AI can improve clinicians’ trust and adoption by providing auditable evidence and interactive interfaces. Unlike prior explainable perception efforts that mainly target medical imaging, DNA structure classification at the molecular level is an upstream keystone in many clinical workflows: predictions must provide not only results (diagnosis/subtype) but also reasons (evidence aligned with clinical concepts).
Despite notable progress of Concept Bottleneck Models (CBMs) in medical imaging (e.g., CEM [1]; AdaCBM [2]; Knowledge-Aligned CBM [3]), three gaps remain: (i) a persistent focus on image modalities with limited, systematic exploration of interpretable modeling and concept definition for sequence data such as DNA; (ii) a lack of end-to-end integration with medical robotic systems, hindering plug-and-play perception–decision coupling in real workflows; and (iii) an “explanation-as-endpoint” paradigm with no closed loop that turns concept evidence into actionable recommendations and next steps, which limits trustworthy clinician-in-the-loop decisions. We propose an automated genomic interpretation system based on a concept bottleneck architecture. The system automates the full pipeline from raw DNA sequences to interpretable concepts, classification, and cost-aware recommendations, reducing manual analysis and serving as a decision module for clinical robotics. Technically, sequences are mapped to images via Chaos Game Representation (CGR), encoded by a CNN, and constrained through a strict Concept Bottleneck Module (CBM). Additional regularizers—concept fidelity, prior alignment, KL matching, and uncertainty calibration—ensure both accuracy and interpretability. A recommendation layer then translates predictions into actionable decisions, making the system transparent, robust, and ready for deployment in medical automation.
@@
Overall, our contributions are summarized as follows: 1. We are the first to combine Chaos Game Representation (CGR) with CBMs for DNA discrimination, yielding a molecular-level, interpretable perception framework that can be seamlessly embedded into medical-robot pipelines—demonstrating potential for human–robot collaborative diagnosis. 2. We propose a joint optimization strategy that simultaneously minimizes task classification loss and concept-consistency loss, preserving accuracy while improving the reliability and interpretability of concept predictions. 3. We design an explanation interface that translates the model’s decision process into clinician-understandable evidence, enabling robots to “explain to humans” and increasing clinical adoptability and transparency in collaboration. 4. We introduce a Recommendation Layer that maps explanations into actionable next-step suggestions, closing the loop from explainable perception to executable recommendations. We further provide a system-level integration case to highlight robotic relevance. Our framework is shown in Fig. 1.
@@ II-A Concept Bottleneck Models
[4] first proposed Concept Bottleneck Models (CBMs), which insert discrete “concept slots” into neural networks to strengthen semantic interpretability. Since then, CBMs have been extended along multiple dimensions: Concept Embedding Models replace discrete slots with continuous embedding spaces to improve robustness under scarce annotations [5]; Stochastic CBMs model processed each concept as a learnable distribution and characterized inter-concept covariances, enabling interventions on one concept to propagate to others [6]. In medical imaging, CBMs with “visual concept filtering” refine the concept set by pruning task-irrelevant visual concepts, thereby further improving the interpretability of diagnostic models [7]. Meanwhile, other studies inject clinical knowledge into black-box models: [8] proposed a regularization module that guides attention maps toward clinically interpretable features in histopathological analysis (e.g., nuclei, lipid droplets); [9] designed a selection module that mimics dermatologists’ prioritization of lesion features such as plaque and scale amid noisy backgrounds; and [10] introduced a multi-attribute attention network to guide the model in learning clinically relevant concepts (e.g., calcification, nodule shape) for predicting thyroid nodule malignancy.
Compared with existing concept-bottleneck or attention-guided studies, our approach differs in two fundamental ways. First, prior work largely focuses on medical imaging modalities and centers on the interpretability of image regions; in contrast, we extend interpretable concepts from visual appearance to the DNA sequence/molecular level, constructing the concept space with biological priors as anchors to capture mechanistic factors across levels and modalities. Second, whereas many existing methods adopt an “interpretation-as-endpoint” paradigm and lack a closed loop to downstream actions, we couple interpretable concepts with robotic decision-making and recommendation, enabling concept interventions to directly influence policies and suggestions rather than remaining at post-hoc explanation.
@@ II-B DNA representation learning
Traditional DNA sequence classification typically relies on extensive manual and expert annotation to determine sequence origin, function, and type; however, in the absence of a definitive “ground truth,” the stability of taxonomic labels is often questioned [11, 12]. Meanwhile, most classical approaches follow an alignment-based paradigm, which entails high time complexity [13] and often depends on additional information such as homology [14], making them ill-suited for very large or evolutionarily divergent datasets. Against this backdrop, Chaos Game Representation (CGR), proposed by Jeffrey in 1990, maps one-dimensional DNA sequences onto a two-dimensional plane via chaotic iteration [15]. The resulting numerical/image representations can serve as genomic signatures that distinguish closely from distantly related organisms [16], consistent with the definition by Karlin and Burge [17]. Using various distance metrics (e.g., Euclidean), CGRs enable alignment-free comparisons and phylogenetic analyses that reveal evolutionary relationships within a cohort, and CGR has thus been regarded as a milestone in graphical bioinformatics [18]. Its quantified variant, Frequency CGR (FCGR), partitions the plane at resolution k into a $2^{k}\times 2^{k}$ grid and tallies k-mer frequencies, yielding a fixed-size numerical matrix/grayscale image. This representation both compresses variable-length sequences and facilitates cross-species signature analysis. FCGR has demonstrated strong scalability across numerous alignment-free applications, overcoming the runtime and scale bottlenecks of alignment-based methods [19], and is often combined with digital signal processing [20] and machine learning methods [21]. Furthermore, FCGR’s fixed-size property naturally suits convolutional neural networks (CNNs) and other image models [22]; [23] achieved 87% accuracy with a simple CNN on 660 DNA sequences across 11 genomic datasets; [24] combined FCGR with ResNet50 to classify SARS-CoV-2 sequences into 11 clades with 96.29% accuracy, outperforming the random-forest-based tool Covidex; and [25] proposed a CGR-based hybrid pipeline that integrates AlexNet, Lasso, and KNN to detect human coronaviruses using 7,951 whole and partial genomes.
@@
Existing studies have clear limitations. Most CBM-based biomedical approaches stop at post-hoc visualization, lacking a direct link from intermediate evidence to actionable outcomes; genomic sequence models often prioritize classification accuracy or depend on complex alignment pipelines, which hinders scalability and precludes real-time use. These factors fall short of the transparency and immediate decision-making required in medical robotics. To address this, we propose an automated genomic interpretation module that integrates CGR preprocessing, a strict concept bottleneck, and regularizers for fidelity, prior consistency, and distribution alignment, ensuring biologically grounded reasoning while preserving discriminative power. In addition, uncertainty calibration and a cost-aware recommendation layer translate concept-level evidence into clinically actionable results. Overall, this design moves genomic analysis beyond black-box prediction toward an interpretable, automated decision system, laying the groundwork for medical automation and robotic integration.
@@ III-B Chaos Game Representation (CGR) Preprocessing
Within the proposed automated genomic interpretation module, Chaos Game Representation (CGR) functions as a deterministic preprocessing stage that converts symbolic DNA sequences into fixed-dimensional, image-like embeddings. This representation is alignment-free and readily consumable by convolutional neural networks (CNNs), providing a standardized interface between raw genomic data and the downstream learning pipeline.
@@ III-D Concept Supervision and Regularization
$c_{\mathrm{GC}}=\frac{n_{\mathrm{G}}+n_{\mathrm{C}}}{L},\quad c_{\mathrm{CpG}}=\frac{n_{\mathrm{CG}}}{L-1},\quad c^{(k)}_{i}=\frac{n^{(k)}_{i}}{L-k+1}.$ (6)
@@ III-E Prior Consistency and Distribution Matching
$\mathcal{R}_{\mathrm{align}}=\sum_{i\in\mathcal{M}^{+}}\!\mathrm{ReLU}\!\big(-(W_{y})_{y^{*},i}\big)+\sum_{i\in\mathcal{M}^{-}}\!\mathrm{ReLU}\!\big((W_{y})_{y^{*},i}\big)\vskip-2.84526pt$ (8)
@@
$\mathcal{R}_{\mathrm{KL}}=D_{\mathrm{KL}}\!\big(p_{\psi}(\cdot\mid\hat{\mathbf{c}})\ \|\ p_{\xi}(\cdot\mid z)\big),$ (10)
@@ III-F Uncertainty Estimation and Calibration
For the automated genomic interpretation module to be deployed in medical automation systems, it is essential that predictions are not only accurate but also trustworthy. To this end, we incorporate explicit mechanisms for uncertainty estimation and calibration, ensuring that downstream recommendations reflect both predictive performance and confidence. We first quantify uncertainty using predictive entropy and assess calibration with the expected calibration error (ECE):
@@
where $M$ is the number of classes, $B$ the number of bins, $n_{b}$ the number of samples in bin $b$, and $\mathrm{acc}(b)$ and $\mathrm{conf}(b)$ its empirical accuracy and mean confidence. Entropy grows with distributional spread, while ECE summarizes the calibration gap across bins. To further improve calibration, we apply temperature scaling to the logits $\mathbf{o}$ using a held-out validation set $\mathcal{V}$:
$\hat{\mathbf{y}}^{(T)}=\mathrm{softmax}(\mathbf{o}/T),$ (13)
@@ III-H Joint Objective and Curriculum Schedule
where $t$ is the training epoch, $T_{w}$ the warm-up length, $T_{r}$ the ramp duration, and $\lambda_{\cdot}^{\max}$ the peak values of each coefficient. Auxiliary losses are thus delayed until the encoder and classifier are stable, and then increased linearly, preventing premature collapse of the strict bottleneck. This design ensures that the system does not merely optimize for accuracy, but jointly enforces interpretability, domain alignment, and decision readiness. By coordinating these objectives under a curriculum schedule, the automated genomic interpretation module achieves both robust convergence and reliable performance—key requirements for its eventual integration into medical automation and robotic systems.
In summary, our methodology integrates CGR-based preprocessing, a concept bottleneck, and regularized training into a coherent framework. Sequences are transformed into 2D representations, encoded by a CNN, and constrained through interpretable concepts with auxiliary regularizers for fidelity, consistency, and calibration. A cost-aware recommendation layer then translates predictions into clinically actionable decisions. This end-to-end design ensures both accuracy and transparent reasoning, providing a reliable foundation for integration into medical automation and robotic systems.
@@ IV-A Datasets
In-house dataset. We first constructed a dataset consisting of gag gene sequences from both subtype B and non-B HIV strains. Specifically, it contains 1,823 subtype B sequences and 1,807 sequences from subtypes A–G. All sequences were preprocessed by removing ambiguous nucleotides, standardizing bases to uppercase (A/C/G/T), and verifying length consistency. On average, each sequence is about 2046 bp long, which ensures comparability across subtypes and provides a controlled setting for classification and concept prediction tasks within our automated pipeline.
@@ IV-B Classification Performance
We first evaluate the automated genomic interpretation module in terms of classification performance, benchmarking it against six representative classifiers on two datasets: (1) our in-house gag dataset and (2) the LANL HIV dataset. The baselines include XGBoost, KNN, SVM, CNN, LASSO, and Logistic Regression.
[…]
@@
Across both datasets, three consistent findings emerge: (1) Traditional baselines such as KNN provide reasonable but clearly inferior performance, highlighting the difficulty of the task. (2) Strong baselines—XGBoost, CNN, Logistic Regression, and LASSO—achieve accuracy $\geq$0.97 and AUROC $\geq$0.88, showing that both tree-based ensembles and deep neural networks are capable of extracting relevant sequence features. (3) Our module surpasses these baselines by further improving F1 and AUROC while maintaining top-level accuracy. Importantly, these gains are achieved through an interpretable concept bottleneck, confirming that accuracy and interpretability are not mutually exclusive.
Since all models share identical splits, hyperparameters, and comparable parameter counts, improvements can be attributed to the module design rather than implementation bias. These results demonstrate that the automated genomic interpretation module achieves state-of-the-art classification while retaining an interpretable reasoning process.
@@ IV-C Concept Prediction Quality
We benchmark against Vanilla-CBM [4], Post-hoc Regressor [27], Clinical-knowledge CBM[3], and AdaCBM[2]. Our module integrates concept loss, prior consistency, ranking regularization, and curriculum scheduling. Evaluation is based on $R^{2}$ (explained variance), Pearson $r$ (linear correlation), and per-concept AUROC (treating concepts as discriminative signals).
@@
The results in Table III highlight clear differences between models. Vanilla-CBM and Post-hoc regressors fail to recover stable biological signals: for CpG and CCC, $R^{2}$ values are negative or near zero, meaning predictions are worse than simply predicting the mean, and correlations plateau around $r\sim 0.65$–0.70. This indicates that naive bottlenecks cannot reconstruct subtle biological variation. Clinical-knowledge CBM and AdaCBM introduce meaningful improvements, especially for GC ($R^{2}=0.702,0.785$) and CpG ($R^{2}=0.165,0.350$), demonstrating the value of expert priors and adaptive mechanisms. However, both still lag significantly in capturing CpG, a notoriously noisy yet clinically relevant signal. By contrast, our automated genomic interpretation module achieves the most faithful recovery across all metrics. For GC, $R^{2}$ reaches 0.874 with correlation $r=0.940$ and AUROC $0.984$, essentially saturating the task. More importantly, on CpG, our method lifts $R^{2}$ from 0.350 (AdaCBM) to 0.615—a relative improvement of over 75%—while correlation rises to 0.846 and AUROC to 0.912, showing robust capacity to model difficult regulatory properties. Similarly, CCC recovery improves from $R^{2}=0.340$ and $r=0.769$ (AdaCBM) to $R^{2}=0.592$ and $r=0.842$, confirming that motif-level structure is faithfully reconstructed. These improvements are consistent across regression and classification metrics, underscoring that the recovered concepts are not only numerically accurate but also operationally discriminative for downstream tasks.
Qualitative inspection further validates these trends: motifs such as CCC, ACG, and blk2 achieve AUROC above 0.95, GC and CCC predictions show clear separation between B and non-B subtypes, and CpG predictions, while still more challenging, are substantially improved. Taken together, these results demonstrate that the automated genomic interpretation module produces interpretable intermediate features with both statistical fidelity and clinical discriminability. Such reliable concept recovery is essential for embedding genomic analysis into medical automation and robotic systems, where predictive accuracy must be coupled with auditable intermediate reasoning to support safe and trustworthy decision-making.
IV-D Faithfulness of the Automated Genetic Interpretation Module
Faithfulness is crucial for automated genomic interpretation, as predictive concepts must genuinely drive decisions rather than serve as superficial correlates. We evaluate this property through two tests: sufficiency, measuring whether predicted concepts alone can sustain classification accuracy, and necessity, quantifying the accuracy drop when a key concept is removed.

### v2

@@ I Introduction
As medical robots are increasingly deployed for triage, specimen handling, and intra-/post-operative decision support, their perception and decision modules often rely on end-to-end deep learning, whose black-box nature undermines trust in safety-critical settings. Recent embodied intelligence and robot-assistant studies have increasingly emphasized structured reasoning, hierarchical cognition, and interpretable intermediate processes for reliable decision-making in complex environments [15, 17, 16, 33, 32, 29]. In parallel, explainable AI has shown strong potential to improve clinical trust by providing interpretable and auditable evidence. While most existing explainable perception studies focus on medical imaging, recent advances in molecular and biomedical learning have also highlighted the importance of structure-aware modeling for biological data [35, 34]. DNA structure classification is an upstream task in many clinical workflows, where predictions must provide not only diagnostic results but also biologically meaningful explanations. Concept Bottleneck Models (CBMs) have demonstrated promising interpretability in medical imaging (e.g., CEM [4], Knowledge-Aligned CBM [22],[18]), yet three gaps remain: (i) limited exploration of interpretable modeling for sequence data such as DNA, (ii) insufficient integration with medical robotic pipelines, and (iii) explanations often treated as endpoints rather than being linked to actionable recommendations for clinician-in-the-loop decision making. We propose an automated genomic interpretation system based on a concept bottleneck architecture that maps raw DNA sequences to interpretable concepts, classification outcomes, and cost-aware recommendations, enabling automated decision support for clinical robotics. Sequences are first transformed into images via Chaos Game Representation (CGR), encoded by a CNN, and constrained through a CBM. Additional regularizers—including concept fidelity, prior alignment, KL matching, and uncertainty calibration—improve both accuracy and interpretability. A recommendation layer further converts predictions into actionable decisions for medical automation. Although evaluated on HIV subtype classification, the framework is disease-agnostic: because it relies on general sequence statistics rather than disease-specific features, it can be readily applied to other genomic tasks such as SARS-CoV-2 variant classification, bacterial genome typing, and broader pathogen detection.
@@
Overall, our contributions are summarized as follows: 1. We combine CGR with CBMs for DNA discrimination, yielding a molecular-level interpretable perception framework that can be embedded into medical-robot pipelines, demonstrating potential for human–robot collaborative diagnosis. 2. We propose a joint optimization strategy that minimizes classification loss and concept-consistency loss, preserving accuracy while improving the reliability and interpretability of concept predictions. 3. We design an explanation interface that translates the model’s decision process into clinician-understandable evidence, enabling robots to explain their reasoning and improving clinical transparency and adoption. 4. We introduce a Recommendation Layer that maps explanations into actionable suggestions, closing the loop from explainable perception to executable decisions and highlighting the system’s relevance for robotic integration. The overall framework is shown in Fig. 1.
@@ II-A Concept Bottleneck Models
[12] first proposed Concept Bottleneck Models (CBMs), which introduce discrete concept slots into neural networks to enhance interpretability. Since then, CBMs have been extended in several directions. Concept Embedding Models replace discrete concepts with continuous embeddings to improve robustness under scarce annotations [3], while Stochastic CBMs model each concept as a learnable distribution to capture inter-concept dependencies [25]. In medical imaging, visual concept filtering refines the concept set by removing task-irrelevant visual cues [11]. Related studies also inject clinical knowledge into black-box models: attention regularization can guide models toward meaningful structures in histopathology [27], , and multi-attribute attention networks capture diagnostic concepts such as calcification and nodule shape for thyroid nodule analysis [21]. Compared with existing concept-bottleneck and attention-guided approaches, our method differs in two key aspects. First, prior work mainly focuses on medical imaging and interpretable visual regions, whereas we extend concept-based reasoning to DNA sequences by constructing biologically grounded concept spaces at the molecular level. Second, rather than treating explanations as post-hoc outputs, we integrate interpretable concepts with robotic decision-making and recommendation, enabling concept-level interventions to directly influence downstream policies [30].
@@ II-B DNA representation learning
Traditional DNA sequence classification often relies on extensive manual annotation to determine sequence origin, function, and type; however, in the absence of definitive ground truth, the stability of taxonomic labels is often debated [1, 20]. Most methods follow alignment-based paradigms, which incur high computational cost [26] and frequently require additional information such as homology [36], limiting scalability to large or highly divergent datasets. To address these issues, Chaos Game Representation (CGR), introduced by Jeffrey in 1990, maps one-dimensional DNA sequences into two-dimensional numerical representations through chaotic iteration [9]. These representations serve as genomic signatures that distinguish organisms across evolutionary distances [19], consistent with the genomic signature concept of Karlin and Burge [10]. CGR enables alignment-free comparisons and phylogenetic analysis using distance metrics (e.g., Euclidean) and has become a foundational method in graphical bioinformatics [23].
@@
Its extension, Frequency CGR (FCGR), divides the plane into a $2^{k}\times 2^{k}$ grid and counts $k$-mer frequencies, producing a fixed-size matrix or grayscale image representation. This encoding compresses variable-length sequences and facilitates cross-species analysis while scaling well to large datasets [7]. FCGR has been widely integrated with digital signal processing [8] and machine learning techniques [6]. Because FCGR produces fixed-size image representations, it naturally suits convolutional neural networks (CNNs) [14]. For example, [24] achieved 87% accuracy using a CNN on 660 DNA sequences across 11 genomic datasets, [2] combined FCGR with ResNet50 to classify SARS-CoV-2 sequences into 11 clades with 96.29% accuracy, and [5] proposed a CGR-based pipeline integrating AlexNet, Lasso, and KNN to detect human coronaviruses from 7,951 genomes. Existing studies have clear limitations. Most CBM-based biomedical approaches stop at post-hoc visualization, lacking a direct link from intermediate evidence to actionable outcomes; genomic sequence models often prioritize classification accuracy or depend on complex alignment pipelines, which hinders scalability and precludes real-time use. These factors fall short of the transparency and immediate decision-making required in medical robotics. To address this, we propose an automated genomic interpretation module that integrates CGR preprocessing, a strict concept bottleneck, and regularizers for fidelity, prior consistency, and distribution alignment, ensuring biologically grounded reasoning while preserving discriminative power. In addition, uncertainty calibration and a cost-aware recommendation layer translate concept-level evidence into clinically actionable results. Overall, this design moves genomic analysis beyond black-box prediction toward an interpretable, automated decision system, laying the groundwork for medical automation and robotic integration.
@@ III-B Chaos Game Representation (CGR) Preprocessing
Within the proposed automated genomic interpretation module, Chaos Game Representation (CGR) deterministically maps symbolic DNA sequences into fixed-dimensional embeddings for downstream learning [31]. This representation is alignment-free and readily consumable by convolutional neural networks (CNNs), providing a standardized interface between raw genomic data and the downstream learning pipeline.
@@ III-D Concept Supervision and Regularization
$c_{\mathrm{GC}}=\frac{n_{\mathrm{G}}+n_{\mathrm{C}}}{L},c_{\mathrm{CpG}}=\frac{n_{\mathrm{CG}}}{L-1},c^{(k)}_{i}=\frac{n^{(k)}_{i}}{L-k+1}.$ (6)
@@ III-E Prior Consistency and Distribution Matching
$\mathcal{R}_{\mathrm{align}}=\sum_{i\in\mathcal{M}^{+}}\!\mathrm{ReLU}\!\big(-(W_{y})_{y^{*},i}\big)+\sum_{i\in\mathcal{M}^{-}}\!\mathrm{ReLU}\!\big((W_{y})_{y^{*},i}\big)\vskip-5.69054pt$ (8)
@@
$\mathcal{R}_{\mathrm{KL}}=D_{\mathrm{KL}}\!\big(p_{\psi}(\cdot\mid\hat{\mathbf{c}})\ \|\ p_{\xi}(\cdot\mid z)\big),\vskip-5.69054pt$ (10)
@@ III-F Uncertainty Estimation and Calibration
For the automated genomic interpretation module to be deployed in medical automation systems, it is essential that predictions are not only accurate but also trustworthy. We incorporate explicit mechanisms for uncertainty estimation and calibration, ensuring that downstream recommendations reflect both predictive performance and confidence. We first quantify uncertainty using predictive entropy and assess calibration with the expected calibration error (ECE):
@@
where $M$ is the number of classes, $B$ the number of bins, $n_{b}$ the number of samples in bin $b$, and $\mathrm{acc}(b)$ and $\mathrm{conf}(b)$ its empirical accuracy and mean confidence. Entropy grows with distributional spread, while ECE summarizes the calibration gap across bins. To further improve calibration, we apply temperature scaling to the logits $\mathbf{o}$ using a held-out validation set $\mathcal{V}$: $\hat{\mathbf{y}}^{(T)}=\mathrm{softmax}(\mathbf{o}/T),$
@@ III-H Joint Objective and Curriculum Schedule
where $t$ is the training epoch, $T_{w}$ the warm-up length, $T_{r}$ the ramp duration, and $\lambda_{\cdot}^{\max}$ the peak values of each coefficient. Auxiliary losses are thus delayed until the encoder and classifier are stable, and then increased linearly, preventing premature collapse of the strict bottleneck. This design ensures that the system does not merely optimize for accuracy, but jointly enforces interpretability, domain alignment, and decision readiness. By coordinating these objectives under a curriculum schedule, the automated genomic interpretation module achieves both robust convergence and reliable performance—key requirements for its eventual integration into medical automation and robotic systems. In summary, our methodology integrates CGR-based preprocessing, a concept bottleneck, and regularized training into a coherent framework. Sequences are transformed into 2D representations, encoded by a CNN, and constrained through interpretable concepts with auxiliary regularizers for fidelity, consistency, and calibration. A cost-aware recommendation layer then translates predictions into clinically actionable decisions. This end-to-end design ensures both accuracy and transparent reasoning, providing a reliable foundation for integration into medical automation and robotic systems.
@@ IV-A Datasets
In-house dataset. We first constructed a dataset consisting of gag gene sequences from both subtype B and non-B HIV strains. Specifically, it contains 1,823 subtype B sequences and 1,807 sequences from subtypes A–G. All sequences were preprocessed by removing ambiguous nucleotides, standardizing bases to uppercase (A/C/G/T), and verifying length consistency. On average, each sequence is about 2046 bp long, which ensures comparability across subtypes and provides a controlled setting for classification and concept prediction tasks within our automated pipeline. For all experiments, the datasets were divided into training, validation, and test sets with a ratio of 70%/10%/20%. To prevent data leakage, sequences originating from the same patient were assigned exclusively to a single split (i.e., patient-level split), thereby ensuring independence between training and evaluation data.
@@ IV-B Classification Performance
We first evaluate the automated genomic interpretation module in terms of classification performance, benchmarking it against six representative classifiers on two datasets: (1) our in-house gag dataset and (2) the LANL HIV dataset. The baselines include XGBoost, KNN, SVM, CNN, LASSO, and Logistic Regression. All models are trained with identical train/val/test splits, capacity-matched within $\pm$20%, and evaluated using Accuracy, F1 Score, and AUROC. Thresholds are fixed based on validation data. All experiments were repeated three times with different random seeds, and the reported results correspond to the average performance across runs.
@@
As shown in Table I, Table II, across both datasets, three consistent findings emerge: (1) Traditional baselines such as KNN provide reasonable but clearly inferior performance, highlighting the difficulty of the task. (2) Strong baselines—XGBoost, CNN, Logistic Regression, and LASSO—achieve accuracy $\geq$0.97 and AUROC $\geq$0.88, showing that both tree-based ensembles and deep neural networks are capable of extracting relevant sequence features. (3) Our module surpasses these baselines by further improving F1 and AUROC while maintaining top-level accuracy. Importantly, these gains are achieved through an interpretable concept bottleneck, confirming that accuracy and interpretability are not mutually exclusive. Since all models share identical splits, hyperparameters, and comparable parameter counts, improvements can be attributed to the module design rather than implementation bias. These results demonstrate that the automated genomic interpretation module achieves state-of-the-art classification while retaining an interpretable reasoning process.
@@ IV-C Concept Prediction Quality
We benchmark against Vanilla-CBM [12], Post-hoc Regressor [28], Clinical-knowledge CBM[22]. Our module integrates concept loss, prior consistency, ranking regularization, and curriculum scheduling. Evaluation is based on $R^{2}$ (explained variance), Pearson $r$ (linear correlation), and per-concept AUROC (treating concepts as discriminative signals).
@@
The results in Table III highlight clear differences between models. Vanilla-CBM and Post-hoc regressors fail to recover stable biological signals: for CpG and CCC, $R^{2}$ values are negative or near zero, meaning predictions are worse than simply predicting the mean, and correlations plateau around $r\sim 0.65$–0.70. This indicates that naive bottlenecks cannot reconstruct subtle biological variation. Clinical-knowledge CBM and AdaCBM introduce meaningful improvements, especially for GC ($R^{2}=0.702,0.785$) and CpG ($R^{2}=0.165,0.350$), demonstrating the value of expert priors and adaptive mechanisms. However, both still lag significantly in capturing CpG, a notoriously noisy yet clinically relevant signal. By contrast, our automated genomic interpretation module achieves the most faithful recovery across all metrics. For GC, $R^{2}$ reaches 0.874 with correlation $r=0.940$ and AUROC $0.984$, essentially saturating the task. More importantly, on CpG, our method lifts $R^{2}$ from 0.350 (AdaCBM) to 0.615—a relative improvement of over 75%—while correlation rises to 0.846 and AUROC to 0.912, showing robust capacity to model difficult regulatory properties. Similarly, CCC recovery improves from $R^{2}=0.340$ and $r=0.769$ (AdaCBM) to $R^{2}=0.592$ and $r=0.842$, confirming that motif-level structure is faithfully reconstructed. These improvements are consistent across regression and classification metrics, underscoring that the recovered concepts are not only numerically accurate but also operationally discriminative for downstream tasks. Qualitative inspection further validates these trends: motifs such as CCC, ACG, and blk2 achieve AUROC above 0.95, GC and CCC predictions show clear separation between B and non-B subtypes, and CpG predictions, while still more challenging, are substantially improved. Taken together, these results demonstrate that the automated genomic interpretation module produces interpretable intermediate features with both statistical fidelity and clinical discriminability. Such reliable concept recovery is essential for embedding genomic analysis into medical automation and robotic systems, where predictive accuracy must be coupled with auditable intermediate reasoning to support safe and trustworthy decision-making.
IV-D Faithfulness of the Automated Genetic Interpretation
Faithfulness is crucial for automated genomic interpretation, as predictive concepts must genuinely drive decisions rather than serve as superficial correlates. As summarized in Table IV, We evaluate this property through two tests: sufficiency, measuring whether predicted concepts alone can sustain classification accuracy, and necessity, quantifying the accuracy drop when a key concept is removed.